### v1

# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

Abstract
Hallucinations in multimodal large language model are caused by the text-visual bias and the co-occurrence bias. The former reflects an over-reliance on text information in the decision-making process, while the latter arises from the statistical object-pairing patterns abstracted from the training data. Existing mitigation methods heuristically address these biases without understanding the fluctuating bias level across the instances. We first propose estimating the influence of respective token types (visual, prompt, and previous outputs) using a gradient-based self-reflection method. The estimated token influence further enables the detection of object-related visual tokens and their integration into an influence-aware contrastive decoding framework to mitigate both types of biases simultaneously. Our method operates without the need for additional resources, such as costly fine-tuning, extra models, or data statistics. Extensive experiments show it effectively reduces hallucinations, achieving up to a $92\%$ accuracy increase on LLaVA-QA90.
1 Introduction
Figure 1: Overview of our method, which acts as a token-wise post-processing mechanism to mitigate text-visual hallucinations by enhancing visual influence (colored bars) in prediction. Our token-influence approach explicitly aligns this amplification with the most influential text inputs—either prompts (gray bar) or previous outputs (white bar). By leveraging token-wise influence measurement, the method also mitigates co-occurrence bias by selectively avoiding the enhancement of visual tokens (green and magenta bars) linked to previously mentioned objects.
Multimodal Large Language Models (MLLMs) liu2024llava ; dai2023instructblipgeneralpurposevisionlanguagemodels ; ye2024mplug ; bai2023qwen ; chen2024internvl , have recently gained significant momentum in generating coherent, context-aware content across diverse domains. Despite their impressive capabilities, these models face a significant challenge: hallucination, where the generated text content is not grounded in the visual inputs rohrbach2018object ; li2023evaluating . This issue undermines the trust in the outputs of MLLMs.
Hallucinations in MLLMs are found to arise from two key biases studied in the literature li2023evaluating ; kang2023impact ; kim2024discovering : text-visual bias, and co-occurrence bias. Text-visual bias arises when a model relies too heavily on textual input, and neglects the visual information. This bias is also linked to the generated text length zhou2023analyzing ; favero2024multi . Longer text increases the risk of hallucinations, often stemming from early predictions that overly influence subsequent content generation. Co-occurrence bias arises from statistical correlations in the training data li2023evaluating , causing models to incorrectly predict objects based on their frequent pairings even when those objects are absent from the visual input. This bias is challenging to address, and currently lacks mitigation solutions that are free from statistical awareness.
Inference-based hallucination mitigation methods leng2024mitigating ; chen2024halc receive more attention than training-based approaches chen2023mitigating ; jiang2024hallucination ; yue2024less ; ben2023mocha ; sun2023aligning ; kang2023impact due to their cost effectiveness, as they eliminate the necessity for data collection, statistical analysis, and model training. Existing inference-related methods leng2024mitigating ; park2024convis ; favero2024multi ; woo2024don mitigate hallucination heuristically but lack a clear understanding of bias severity. Among inference-related methods, revision approaches xing2024efuf ; deng2024seeing ; radford2021learning often require additional models for prediction validation, such as MLLM-based scoring, which may also hallucinate. Meanwhile, image-level methods leng2024mitigating ; favero2024multi ; zhao2024mitigating fail to address co-occurrence hallucinations at the token (individual object) level. Existing token-level approaches woo2024don ; kan2024catchcomplementaryadaptivetokenlevel ; chen2024halc either lack semantic awareness of visual tokens or rely on auxiliary segmentation or detection models to obtain decoupled visual representations. Therefore, a method that jointly and quantitatively addresses all biases without requiring additional data, or models remains a challenging and unresolved issue.
In this work, to quantitatively estimate biases, we introduce a gradient-based token influence, which leverages the first-order Taylor expansion spivak1980calculus to estimate how each input token influences the output. This token influence enables the detection of the corresponding visual token of each predicted object (noun). In addition, we incorporate both token influence and the detected object-related visual tokens into a contrastive decoding framework li2022contrastive and propose Gradient-based Influence-Aware Contrastive Decoding (GACD) to jointly mitigate text-visual and co-occurrence biases. To address text-visual bias, GACD enhances visual token influence, aligning it with the most influential text components from prompts or previous outputs. This ensures visual grounding answer and prevents visual forgetting. To mitigate the co-occurrence bias, GACD selectively avoids enhancing visual tokens linked to objects mentioned in previous outputs. As illustrated in Fig. 1, GACD corrects the hallucinated prediction of ‘beer’ by enhancing the influence of visual tokens unrelated to ‘fork’ or ‘spoon’, resulting in a more visually grounded output. As text-visual bias tends to occur in long sentence generation zhou2023analyzing , we further introduce a sample-dependent stopping criterion based on visual influence to explicitly capture visual contributions and account for length effects.
Contributions of this work is summarized as:
  • We propose GACD, a novel hallucination mitigation method that reduces text-visual bias by enhancing the influence of visual tokens, without requiring external models or fine-tuning. Its token-wise adjustment further enables the mitigation of co-occurrence bias without relying on training dataset statistics.
  • We propose token influence to express MLLM biases, providing a detailed understanding at both the token (individual object) level and the sample level.
  • Extensive experiments demonstrate that GACD effectively mitigates hallucinations and enhances accuracy without sacrificing information. GACD achieves up to $8\%$ increase in overall score on AMBER wang2023llm , an $8\%$ F1 boost on POPE li2023evaluating , up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 liu2024visual .
2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data mckenna2023sources ; huang2025survey , while in MLLMs, studies tonmoy2024comprehensive ; li2023evaluating ; fu2024mmecomprehensiveevaluationbenchmark show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in favero2024multi ; zhou2023analyzing . Existing methods li2023evaluating ; fu2024mmecomprehensiveevaluationbenchmark ; kim2024discovering typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  kim2019learning ; guo2024bias , and enabling self-reflective hallucination mitigation.
Hallucination Mitigation. Training-related hallucination mitigation methods chen2023minigpt ; jiang2024hallucination ; yue2024less ; ben2023mocha ; sun2023aligning ; lee-etal-2022-deduplicating are expensive, requiring access to training data and specialized statistical analysis to effectively mitigate hallucinations. Reinforcement-learning approaches martino2023knowledge ; xing2024efuf ; deng2024seeing ; radford2021learning ; zhai2024hallecontrolcontrollingobjecthallucination rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text-only LLMs, such methods aim to align outputs with factual knowledge chuang2023dola ; lee2022factuality ; li2023inference . In MLLMs, post-decoding strategies emphasize the role of visual inputs li2023inference ; leng2024mitigating ; zhao2024mitigating ; favero2024multi and can be classified into image-level and token-level interventions. Image-level decoding methods leng2024mitigating ; zhao2024mitigating ; favero2024multi ; zhang2024debiasing ; manevich2024mitigatinghallucinationslargevisionlanguage treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation chen2024halc and detection models kan2024catchcomplementaryadaptivetokenlevel or lack awareness of object-related decoupling woo2024don . In contrast, our method estimates embedded bias and decouples object-aware visual tokens, enabling comprehensive and sample-specific hallucination mitigation without relying on external data or models.
3 Method
In this section, we provide background on MLLMs, introduce the concept of token influence, and explain how GACD balances token influence through contrastive decoding to mitigate hallucinations.
3.1 Background on MLLMs
Consider a sequence $\mathbf{t}^{p}=[t^{p}_{1},\dots,t^{p}_{N}]$ as the input prompt, where $t^{p}_{n}$ ($1\leq n\leq N$) is a prompt token from a predefined vocabulary. Visual tokens $\mathbf{t}^{v}=[t^{v}_{1},\dots,t^{v}_{S}]$ are extracted from visual inputs $V$. The MLLMs generate the response sequence $\mathbf{y}=[y_{1},\dots,y_{M}]$ using the logit generation function $\mathcal{F}_{\theta}(\cdot)$ and the softmax function $\sigma(\cdot)$ 111 Probability of softmax output is interpreted as confidence. as:
$\mathbf{y}=\sigma(\mathcal{F}_{\theta}(\mathbf{t}^{v},\mathbf{t}^{p})).$ (1)
$y_{m}$ denotes an individual output token for $1\leq m\leq M$. The conditional probability distribution $p(\mathbf{y}|\mathcal{F}_{\theta},\mathbf{t}^{p},\mathbf{t}^{v})$ can therefore be expressed as:
$p(\mathbf{y}|\mathcal{F}_{\theta},\mathbf{t}^{v},\mathbf{t}^{p})=\prod_{m=1}^{M}p(y_{m}|\mathcal{F}_{\theta},\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m}),$ (2)
where $\mathbf{y}_{<m}=[y_{1},\dots,y_{m-1}]$ for $m>1$ and is empty for $m=1$. Pre-trained MLLMs encode not only factual knowledge but also statistical correlations and biases from the training data kim2019learning ; guo2024bias . These biases are encoded within the model parameters ($\theta^{\star}$), enabling self-reflective bias interpretation.
3.2 Gradient-Based Token Influence Estimation
To capture these learned biases, we measure the impact each individual token has on the output logits. To achieve this, we propose token influence estimation based on a first-order Taylor expansion spivak1980calculus on logit generation function $\mathcal{F}_{\theta^{*}}$  [^2] Taylor expansion details are in supplementary Sec. A.:
$\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}\approx\sum^{S}_{s=1}\mathbf{g}^{v}_{ms}{t}^{v}_{s}+\sum^{N}_{n=1}\mathbf{g}^{p}_{mn}{t}^{p}_{n}+\sum^{m-1}_{i=1}\mathbf{g}^{y}_{mi}y_{i}+Const,$ (3)
where
$\mathbf{g}^{v}_{ms}:=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{v}_{s}}\bigg{|}_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}},\quad\mathbf{g}^{p}_{mn}:=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{p}_{n}}\bigg{|}_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}},\quad\mathbf{g}^{y}_{mi}:=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial y_{i}},$ (4)
are the first order gradients w.r.t., visual tokens $\mathbf{t}^{v}$, text prompt $\mathbf{t}^{p}$ and previous outputs $y_{i}$ and $Const$ denotes all other terms that are constant w.r.t., the $\mathbf{t}^{v},\mathbf{t}^{p}$.
We measure the importance of input tokens using the norm of their gradients. Specifically, we use the Manhattan (L1) norm, as it effectively captures the contribution of individual tokens. Accordingly, we represent the influence of each token as $\big{|}\mathbf{g}^{v/p/y}_{m\cdot}\big{|}$. The overall influence of visual tokens, text prompts, and previous outputs can be represented as:
$\texttt{I}^{v}_{m}=\sum^{S}_{s=1}\big{|}\mathbf{g}^{v}_{ms}\big{|},\quad\texttt{I}^{p}_{m}=\sum^{N}_{n=1}\big{|}\mathbf{g}^{p}_{mn}\big{|},\quad\texttt{I}^{y}_{m}=\sum^{m-1}_{i=1}\big{|}\mathbf{g}^{y}_{mi}\big{|}.$ (5)
This expression allows us to interpret bias at the individual sample level. We also define the influence of each visual token on a specific output vocabulary token $c$ as:
$I_{ms}(c)=\big{|}\mathbf{g}^{v}_{ms}[c]\big{|},$ (6)
where $\mathbf{g}^{v}_{ms}[c]$ represents the gradient from the output vocabulary $c$ with respect to the input tokens.
3.3 Influence-Aware Contrastive Decoding
We extend contrastive decoding li2022contrastive by incorporating object-aware negative guidance and dynamically computing decoding weights based on token influence. As shown in Fig. 2, for each $m^{th}$ output, contrastive logits are generated alongside the original logits, with the decoding weight $\alpha_{m}$ inferred based on token influence from both.
Figure 2: Overview of the proposed GACD method. GACD adjusts contrastive decoding based on token influence (Sec. 3.2), explicitly amplifying the influence of visual tokens $\mathbf{t}^{u}$. The contrastive logits, acting as negative guidance, are generated by object-related visual token $\mathbf{t}^{o}$, prompt token $\mathbf{t}^{p}$, and previous output token $\mathbf{t}^{y}$, with object-related visual token $\mathbf{t}^{o}$ included only for noun predictions to reduce the influence of potential anchor objects and mitigate co-occurrence bias. The decoding weight $\alpha_{m}$ is calculated based on token influence in both original and contrastive logits, amplifying object-unrelated visual token $\hat{\mathtt{I}}^{u}_{m}$ to align with the dominant text component $\max({\color[rgb]{0,0,1}\hat{\mathtt{I}}^{p}_{m}},{\color[rgb]{0.0,0.5,0.0}\hat{\mathtt{I}}^{y}_{m}})$. For noun predictions, GACD detects the highest-influence visual token using individual visual influence $I_{ms}(y_{m})$ and flags it as object-related $\mathbf{t}^{o}$ for subsequent processing.
We begin with object-related visual token detection. To identify objects in previous outputs $\mathbf{y}_{<m}$, we first perform noun detection, treating each noun as an object. For each identified object, a mask is applied to flag the highest-influence object-related visual tokens. The mask $\mathcal{M}_{ms}$ fuses visual tokens related to all previously mentioned objects:
[…]
where $\bigvee$ represents the bitwise OR operator. The visual tokens are categorized into two groups: object-related visual tokens $\mathbf{t}^{o}$, and Unrelated Object (UO) tokens $\mathbf{t}^{u}$, defined as follows:
$\mathbf{t}^{o}=\mathbf{t}^{v}\odot\mathcal{M}_{m},\quad\mathbf{t}^{u}=\mathbf{t}^{v}\odot(1-\mathcal{M}_{m}),$ (8)
where $\odot$ indicates Hadamard product. The overall influence of these groups is defined as:
$\texttt{I}^{o}_{m}=\sum^{S}_{s=1}(\big{|}\mathbf{g}^{v}_{ms}\big{|}\odot\mathcal{M}_{ms}),\quad\texttt{I}^{u}_{m}=\sum^{S}_{s=1}(\big{|}\mathbf{g}^{v}_{ms}\big{|}\odot(1-\mathcal{M}_{ms})),$ (9)
In noun predictions, this mask removes the impact of tokens related to previously mentioned objects $\mathbf{t}^{o}$ to mitigate co-occurrence bias. For non-noun predictions, this mask calculation is not triggered; all elements in $\mathcal{M}_{m}$ are set to 0, resulting in an empty $\mathbf{t}^{o}$ set.
To jointly mitigate text-visual and co-occurrence biases, GACD adjusts the logits as follows:
$\hat{\mathcal{F}}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}=(1+\alpha_{m})\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\alpha_{m}\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m},$ (10)
where $\alpha_{m}$ is the contrastive decoding weight. This contrastive decoding aims to enlarge the Kullback-Leibler (KL) divergence  [^3] A detailed explanation of why contrastive decoding enlarges KL divergence is in supplementary Sec. B. between the contrastive logits distribution $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}$ and the original logits distribution $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}$. Here, the original logits distribution can be viewed as $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{u},\mathbf{t}^{o},\mathbf{t}^{p})_{m}$, a joint distribution with additional $\mathbf{t}^{u}$. Thus, increasing the KL divergence effectively amplifies the influence of these UO visual tokens $\mathbf{t}^{u}$.
The contrastive decoding weight, $\alpha_{m}$, controls the KL divergence distance and, from another perspective, regulates the influence of each component on the final GACD logits. The influence of UO visual tokens $\hat{\mathtt{I}}^{u}_{m}$, object-related visual tokens $\hat{\mathtt{I}}^{o}_{m}$, the prompt $\hat{\mathtt{I}}^{p}_{m}$, and previous outputs $\hat{\mathtt{I}}^{y}_{m}$ is determined by their gradient values, based on (9) and (5). Follow the chain rule, UO visual tokens are included only in the original logits $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}$ and are amplified by $(1+\alpha_{m})$, while other inputs contribute to the additional contrastive logit $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}$, experiencing smaller gradient changes.
To mitigate text-visual bias, GACD balances the influence of UO visual tokens to align with the dominant text component, defined as $\texttt{I}^{t}_{m}=\max(\texttt{I}^{p}_{m},\texttt{I}^{y}_{m})$. Aligning UO visual token influence with the question prompt $\texttt{I}^{p}_{m}$ is crucial for visually grounded responses, while balancing it with previous outputs $\texttt{I}^{y}_{m}$ is essential for preventing visual forgetting. Under this condition, $\alpha_{m}$ is calculated as:
[…]
where $\tilde{\mathtt{I}}^{o}_{m}$, $\tilde{\mathtt{I}}^{p}_{m}$, and $\tilde{\mathtt{I}}^{y}_{m}$ represent the influence of object-related visual tokens, prompt, and previous output tokens, respectively, on the contrastive logits $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}$, calculated follow (5).
Furthermore, unlike existing contrastive decoding methods zhou2023analyzing ; favero2024multi ; leng2024mitigating , which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our method explicitly imposes a constraint to prevent the influence of object-related visual and prompt from becoming negative:
$\alpha_{m}=\min(\alpha_{m},\frac{\mathtt{I}^{o}_{m}}{\tilde{\mathtt{I}}^{o}_{m}-\mathtt{I}^{o}_{m}},\frac{\mathtt{I}^{p}_{m}}{\tilde{\mathtt{I}}^{p}_{m}-\mathtt{I}^{p}_{m}}).$ (12)
Additionally, to further mitigate text-visual bias in long-term generation, we propose a sample-dependent stopping criterion based on the visual influence. If the visual influence ratio $r^{v}_{m}$ for the token immediately following the end-of-sequence (EOS) token falls below a defined threshold $\epsilon$:
$r^{v}_{m}=\frac{\mathtt{I}^{v}_{m}}{\mathtt{I}^{v}_{m}+\mathtt{I}^{p}_{m}+\mathtt{I}^{y}_{m}}<\epsilon\text{ and }y_{m-1}=\text{EOS},$ (13)
early stopping is triggered to prevent further output generation with minimal visual grounding.
4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We follow leng2024mitigating ; woo2024don ; deng2024seeing ; huang2024opera to use Amber wang2023llm , MSCOCO lin2014microsoft and LLaVa-QA90 liu2024visual datasets for the generative task, and Amber wang2023llm and POPE li2023evaluating datasets on the discriminative tasks.
[…]
Implementation Details. The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling graves2013generating .
4.1 Results on Open-ended Generation
In this section, we compare against SOTA alignment-based method RLAIF-V yu2024rlaif and contrastive decoding methods VCD leng2024mitigating , M3ID favero2024multi , and AVISC woo2024don , on the AMBER and MSCOCO datasets, as presented in Tab. 1, Tab. 2. Additionally, we evaluate our method against VCD on the LLaVA-QA90 dataset, presented in Tab. 3. Our method outperforms most existing approaches across various baseline models and datasets, highlighting its robustness and reliable performance across different data types and model architectures. Specifically, we surpass image-level contrastive decoding methods like VCD and M3ID, demonstrating its effectiveness in operating at the token level and adapting to individual samples. Furthermore, compared to the token-level AVISC, our method excels, likely due to its object awareness and adaptability to fluctuating bias levels. The results further demonstrate that our method effectively mitigates hallucinations while preserving information.
Table 1: Results on the AMBER Dataset.
MLLMs Method Generative Discriminative Score $\uparrow$
[…]
LLaVA v1.5 base 7.8 51.0 36.4 4.2 72.0 93.2 62.4 74.7 83.5
RLAIF-V 6.6 49.7 32.0 2.9 76.7 92.0 78.1 84.5 89.0
VCD 6.7 46.4 32.6 3.5 71.3 91.1 62.3 74.3 83.8
M3ID 6.2 50.5 29.3 2.8 72.4 91.8 64.1 75.5 84.7
AVISC 6.5 50.2 34.8 2.7 73.8 89.7 68.4 77.6 85.5
Ours 5.6 51.0 24.3 1.8 80.3 82.9 89.3 86.0 90.2
Instruct BLIP base 8.8 52.2 38.2 4.4 76.5 84.5 79.0 81.7 86.5
RLAIF-V 7.6 47.7 29.9 2.8 76.5 84.5 79.0 81.7 87.1
VCD 7.9 49.7 36.7 3.7 75.9 83.5 79.3 81.3 86.7
M3ID 7.3 49.2 33.8 3.7 75.8 84.4 77.9 81.0 86.9
AVISC 7.1 48.8 34.4 4.3 75.9 83.4 79.5 81.4 87.2
Ours 6.0 49.4 26.6 2.4 78.1 88.8 76.6 82.2 88.1
mPLUG Owl2 base 10.6 52.0 39.9 4.5 75.6 95.0 66.9 78.5 84.0
RLAIF-V 7.8 50.5 35.7 4.1 81.2 90.8 79.7 84.9 88.6
VCD 8.0 51.3 35.3 4.1 75.6 83.5 78.8 81.1 86.6
M3ID 7.8 51.7 34.9 4.1 75.9 83.5 79.3 81.3 86.8
AVISC 10.9 50.5 35.5 4.4 82.1 90.7 81.4 85.8 87.5
Ours 7.5 53.6 34.7 4.0 82.1 87.0 86.2 86.6 89.6
Table 2: Results Using the CHAIR Metric on the MSCOCO Subset Following deng2024seeing .
Method LLaVA-v1.5 InstructBLIP mPLUG-Owl2
$C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$
base 48.8 13.4 78.6 99.8 57.8 16.5 73.6 101.3 59.2 17.6 75.8 105.3
VCD 44.8 12.8 76.8 89.8 63.4 19.6 71.2 95.5 52.7 16.1 73.2 93.6
M3ID 44.5 12.1 77.0 85.1 57.3 16.1 72.5 100.1 52.4 15.8 72.7 92.6
AVISC 46.4 13.4 76.3 90.5 58.9 17.8 70.6 99.6 58.3 17.5 75.6 99.5
Ours 41.0 10.9 77.3 85.0 47.4 13.4 72.3 93.9 45.0 12.4 74.9 83.5
Hallucination Mitigation. Our approach achieves hallucination reductions up to $33\%$ at sentence-level (halhal in Tab. 1 and $C_{S}$ in Tab. 2) and $32\%$ at instance-level (chacha in Tab. 1 and $C_{I}$ in Tab. 2), demonstrating its effectiveness in mitigating overall hallucinations. It also effectively addresses co-occurrence hallucinations, achieving up to a $57\%$ reduction for cogcog in Tab. 1. We detail the text-visual and co-occurrence bias analysis in Sec. 5. Furthermore, an accuracy increase of up to $92\%$ in Tab. 3 underscores our model’s effectiveness in reducing hallucinations by staying closely aligned with the input image. This improvement highlights our method’s capacity to jointly mitigate multiple types of hallucinations.
Information Preservation. Our method also enhances information preservation, with recall (covcov in Tab. 1 and $R$ in Tab. 2) dropping by an average of only $1.1\%$, compared to an average drop of $3.2\%$ in other methods, and even increasing by $3.1\%$ when using the baseline mPLUG-Owl2 on the AMBER dataset. Higher recall indicates that our model retrieves a broader range of objects from visual inputs. Additionally, results in Tab. 3 an increase of up to $45\%$ in detailedness ($Det$), further demonstrating our method’s effectiveness in retrieving all relevant visual details and mitigating visual forgetting.
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following leng2024mitigating .
Method LLaVA-v1.5 IntructBLIP mPLUG-Owl2
[…]
base 3.23 3.54 3.84 4.07 4.07 4.33
VCD 4.15 3.85 4.23 4.69 4.52 4.64
Ours 6.20 5.13 6.28 4.77 6.69 6.28
4.2 Results on Discriminative Task
Next, we evaluate our method on discriminative tasks using AMBER (discriminative VQA) and POPE (existence VQA) datasets, with results shown in Tab. 1 and Tab. 4. The findings confirm that our method achieves a better balance between precision and recall, resulting in consistently higher F1 scores and improved accuracy. Notably, for the high-performance baseline Intern-VL2, other methods degrade its original performance, whereas ours effectively preserves it, likely due to its bias awareness. Additionally, we observe variation in performance improvements across different question categories and baseline MLLMs.
Table 4: Results on POPE in MSCOCO Adversarial Setting.
Method LLaVA-v1.5 InstructBLIP mPLUG-Owl2 InternVL2
$uparrow Acc\uparrow F1\uparrow$ $uparrow Acc\uparrow F1\uparrow$
base 79.4 81.6 79.8 81.4 72.5 77.5 85.8 85.0
VCD 80.9 81.3 79.6 79.5 74.2 78.8 83.2 82.2
M3ID 81.7 81.8 81.0 81.6 75.6 79.1 83.5 82.1
AVISC 81.2 81.6 81.8 81.9 80.9 79.7 85.3 84.6
Woodpecker 80.5 80.6 79.0 78.6 77.5 76.9 85.7 84.8
Ours 83.5 82.1 82.5 82.1 84.2 83.7 85.8 85.0
(a) Visual Influence Ratio
(b) F1 across Question Categories
Figure 3: (a) Visual influence ratios across the POPE dataset, illustrating variation across MLLMs. Our method successfully increases the visual influence ratio when it falls below $50\%$. (b) F1 scores for the AMBER discriminative task using LLaVA-v1.5 are consistently improved by our method, with particularly notable gains in the existence and state categories.
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using baseline model LLaVA-v1.5 liu2024llava . Our method improves performance across all categories, with particularly significant gains in existence, attributes, and state. These improvements can be attributed to the enhanced influence of visual tokens, which benefit categories that are more readily discernible from visual inputs.
Variation in Improvement Across MLLMs. Our method achieves the most significant improvement on mPLUG-Owl2 in object existence VQA (Tab. 4) and on LLaVA-v1.5 across various VQA categories (Tab. 1 discriminative), with minor improvement observed on InternVL2. This variation in performance correlates with the baseline visual influence ratios of the MLLMs. Fig. 3(a) presents the visual influence ratios in object existence VQA, showing that LLaVA-v1.5 exhibits the lowest visual contribution, followed by mPLUG-Owl2. This lower baseline visual influence ratio allows our method to make more impactful adjustments. In contrast, InternVL2 has an original visual influence ratio exceeding $50\%$, resulting in minimal improvement when our method is applied. The strong performance of InternVL2 can be attributed to its original high visual influence ratio, further validating the motivation behind our approach. However, this dominant visual influence ratio in InternVL2 is not consistent across other question categories or in open-ended generation tasks  [^4] Results for other question categories and visualizations of open-ended generation are in supplementary Tab. 20 and Fig. 9., which allows our method to enhance its performance on the AMBER dataset (Tab. 6).
4.3 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6 and Intern-VL2, on the Amber dataset. The results in Tab. 6 in supplementary demonstrate that our method consistently improves MLLM performance, even when their original performance is high. To further evaluate the effectiveness of our method across diverse tasks, we conduct evaluations on the MMBench liu2025mmbench and MMVet yu2023mm datasets. The results are presented in supplementary Tab. 8 and Tab. 9. The evaluation results show that our method enhances perception and recognition performance, particularly in tasks where visual contribution is straightforward. Moreover, it maintains overall task performance, demonstrating its robustness.
5 Ablation Study
In this section, we first analyze text–visual and co-occurrence biases, and then study the contributions of our proposed components. Further details on gradient computation methods, norm selection, and hyperparameter settings are provided in the supplementary material.
Figure 4: Comparison of visual influence ratios $r^{v}_{m}$ and Text-Visual GAP, with and without our GACD. (a) Without GACD, the baseline mPLUG-Owl2 shows a low initial visual influence ratio, punctuation marks and suffixes naturally have low visual influence, while objects in images initially have relatively high visual influence, which decreases as sequence length grows. Hallucinations tend to occur when the visual ratio is low. The text-visual gap confirms that text dominates the influence on predictions. (b) With GACD, the visual influence ratio increases overall and mitigates the decrease over the sequence length. The text tokens only domain influence in predictions less related to the visual, reducing hallucination.
[…]
(a) Co-occurrence Hallucination
(b) Individual Visual Token Influence
(c) Overall Statistics
Figure 5: Co-occurrence hallucination of ‘table’ in the presence of ‘chair’. (a) Comparison of outputs with and without GACD. (b) Visualization of individual visual token influence indicates that the visual token with index 24, which has the highest influence on the hallucinated ‘table’, also holds the highest influence on ‘chair’. (c) Summary statistics for 100 chair-only and 100 table-only images, showing the hallucination rate and the percentage of cases where both objects share the same most influential visual token (as illustrated in b). GACD effectively reduces both metrics.
Co-occurrence Bias Analysis. Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’ in the image. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset lin2014microsoft . Experimental results in Fig. 5(c) indicate that when either only ‘chair’ or ‘table’ exists in the image, the hallucination rate for the other object is $23.5\%$, with a shared most influential visual token rate of $31.9\%$ in these hallucinations. This indicates that the ‘Same Most Influential Token’ phenomenon is common in co-occurrence hallucinations. Our GACD effectively reduces the hallucination where both ‘table’ or ‘chair’ are predicted in single-object images.
5.1 Component Analysis
To assess the effectiveness of each component in our proposed method, we conducted the following analyses: 1) Visual Amplification (VA) only: Applied visual amplification to all visual tokens, including during noun predictions. 2) Co-occurrence Hallucination Reduction (CR): Detected object-related visual tokens and amplified UO visual tokens during noun predictions. 3) Our full model, with early stopping (ES). Tab. 5 demonstrates that each component of our method contributes to the overall performance. VA significantly reduces hallucinations while improving object recall. CR further mitigates co-occurrence bias, a residual form of the text-visual bias addressed by VA, resulting in additional hallucination reduction. When needed, ES mechanism shortens output length to effectively reduce hallucinations with only a minor trade-off in object recall.
Table 5: Component Analysis Using the CHAIR Metric.
Components LLaVA-v1.5 InstructBLIP mPLUG-Owl2
VA CR ES $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$
48.8 13.4 78.6 99.8 57.8 16.5 73.6 101.3 59.2 17.6 75.8 105.3
46.4 11.6 79.0 95.6 53.6 15.1 75.3 108.4 52.6 14.4 78.2 95.6
46.2 11.3 79.4 95.5 53.2 14.0 74.6 105.7 52.3 14.2 78.0 95.5
41.0 10.9 77.3 85.0 47.4 13.4 72.3 93.9 45.0 12.4 74.9 83.5
6 Conclusion
In conclusion, we introduce a gradient-based self-reflection method to estimate token influence and quantitatively estimate bias severity. This estimation enables the identification of object-related visual tokens, which are then integrated into an influence-aware contrastive decoding framework. This framework effectively mitigates both text-visual and co-occurrence biases, reducing hallucinations. Our method operates without requiring additional resources such as costly fine-tuning, extra models, or data statistics. Furthermore, to reduce text-visual bias in long-generated sequences, we propose a sample-dependent stopping criterion based on the proposed visual influence.
Appendix A First Order Taylor Expansion
The first order Taylor expansion of $\mathcal{F}_{\theta^{*}}$ w.r.t., $\mathbf{t}^{v},\mathbf{t}^{p}$ and, all the previous output tokens $\mathbf{y}_{i}$ with $i<m$:
$\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}\approx$ $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)})_{m}+\sum^{S}_{s=1}\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{v}_{s}}\bigg{|}_{\mathbf{t}^{v(0)}}({t}^{v}_{s}-{t}^{v(0)}_{s})$ (14)
$+\sum^{N}_{n=1}\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{p}_{n}}\bigg{|}_{\mathbf{t}^{p(0)}}({t}^{p}_{n}-{t}^{p(0)}_{n})+\sum^{m-1}_{i=1}\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial y_{i}}\bigg{|}_{y^{(0)}_{i}}(y_{i}-y^{(0)}_{i})$
$=$ $\sum^{S}_{s=1}\mathbf{g}^{v}_{ms}{t}^{v}_{s}+\sum^{N}_{n=1}\mathbf{g}^{p}_{mn}{t}^{p}_{n}+\sum^{m-1}_{i=1}\mathbf{g}^{y}_{mi}y_{i}+Const,$
where $\mathbf{g}^{v}_{ms}=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{v}_{s}}\Big{|}_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}}$, $\mathbf{g}^{p}_{mn}=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{p}_{n}}\Big{|}_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}}$ and $\mathbf{g}^{y}_{mi}=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial y_{i}}\Big{|}_{\mathbf{y}_{i}=\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)})}$ are the first order gradients w.r.t., visual tokens $\mathbf{t}^{v}$, text prompt $\mathbf{t}^{p}$ and previous outputs $y_{i}$ and $Const$ denotes all other terms that are constant w.r.t., the $\mathbf{t}^{v},\mathbf{t}^{p}$.
Appendix B Interpreting Contrastive Decoding through KL Divergence
Kullback-Leibler (KL) divergence can be used to interpret contrastive decoding, For example, in GACD VA mode, it measures the divergence between the reference distribution $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})$ to the visual joint distribution $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})$.
$\mathrm{D}_{KL}$ $=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})\text{log}\big{(}\frac{p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})}{p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})}\big{)}$ (15)
$=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\text{log}(p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m}))-\text{log}(p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})))$
$=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}[c]-\text{log}(\sum\exp(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}))$
$-\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m}[c]+\text{log}(\sum\exp(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m})))$
$=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\underbrace{(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m})}_{\text{adjustment term}}[c]+Const),$
where $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})=\mathbf{\sigma}(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m})$, $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})=\mathbf{\sigma}(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m})$ and $c$ represents a class in the predefined vocabulary. The adjustment term $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m}$ increases the KL divergence, thereby emphasizing the impact of visual tokens.
Appendix C Result on modern MLLMs
We further evaluate modern MLLMs, LLaVA-v1.6 and Intern-VL2, on the Amber dataset. The results in Tab. 6 demonstrate that our method consistently improves MLLM performance, even when their original performance is high.
Table 6: Latest MLLMs on the AMBER Dataset.
MLLMs Method Generative Discriminative Score $\uparrow$
[…]
LLaVA v1.6 base 9.9 56.7 47.4 4.3 80.3 82.9 89.3 86.0 88.5
Ours 8.7 58.3 43.8 2.5 81.2 85.2 88.8 87.0 89.2
Intern VL2 base 8.1 69.6 59.0 5.2 84.0 87.3 88.8 88.0 90.0
Ours 7.9 69.8 57.8 3.7 84.7 88.2 88.8 88.5 90.3
Appendix D MLLMs Architectures
Tab. 7 shows detailed information about the vision encoder and LLM components of the MLLM architectures used in our experiments.
Table 7: Details of the used MLLM architectures.
MLLMs Vision encoder LLM
LLaVA-v1.5 (7B) CLIP-L-336px Vicuna-v1.5-7B
LLaVA-v1.5-13B CLIP-L-336px Vicuna-v1.5-13B
LLaVA-v1.6 CLIP-L-336px Vicuna-v1.5-7B
InstructBLIP (7B) BLIP-2 Vicuna-v1.1-7B
InstructBLIP-13B BLIP-2 Vicuna-v1.1-13B
mPLUG-Owl2 CLIP-L LLaMA-2-7B
InternVL2-4B InternViT-300M-448px Phi-3-mini-128k-instruct
Appendix E Results on MMBench
We further evaluate our method on MMBench [33]. The results in Tab. 8 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
Table 8: Results on MMBench Dataset.
MLLMs Method Overall CP FP-S FP-C AR LR $R$
LLaVA-V1.5 base 62.3 68.5 69.6 57.7 73.1 29.9 54.7
ours 61.8 73.2 62.6 53.0 73.3 27.8 57.8
mPLUG-Owl2 base 63.5 68.1 69.1 55.8 78.4 37.0 57.0
ours 65.0 72.6 66.6 53.0 76.0 41.6 63.0
Appendix F Results on MM-Vet
The evaluation on MM-Vet [49] in Tab. 9 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.
Table 9: Results on MM-Vet dataset.
MLLMs Method Rec OCR Know Gen Spat Math Total
LLaVA-V1.5 base 32.9 20.1 19.0 20.1 25.6 5.2 28.0
ours 38.9 24.9 15.0 15.5 24.9 7.7 28.9
InstructBlip base 32.4 14.6 16.5 18.2 18.6 7.7 26.2
ours 40.5 18.0 18.7 17.4 14.9 3.8 26.6
mPLUG-Owl2 base 36.1 19.4 29.8 19.4 23.9 7.7 27.3
ours 45.0 26.4 27.9 25.9 24.8 3.8 33.9
Appendix G Results on MME
Our evaluation on MME [13] dataset is presented in Tab. 10. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
Table 10: Result on MME Dataset.
MLLMs Method Existence $\uparrow$ Count $\uparrow$ Position $\uparrow$ Color $\uparrow$ Total $\uparrow$
LLaVA-v1.5 base 190.0 140.0 128.3 155.0 613.3
ours 190.0 153.3 128.3 163.3 634.9
IntructBLIP base 180.0 55.0 50.0 130.0 415.0
ours 185.0 55.0 50.0 130.0 420.0
mPLUG-Owl2 base 170.0 145.0 73.3 158.3 546.6
ours 170.0 150.0 73.3 150.0 548.3
Appendix H Other Results of POPE
We report our experimental results on the POPE dataset, in addition to MSCOCO and adversarial settings, in Tab. 11. The results indicate that our method improves performance across all baseline MLLMs, with more significant gains observed in the adversarial setting. This discrepancy likely arises because adversarial scenarios require models to rely more heavily on visual inputs, aligning with our method’s focus on enhancing visual influence. Conversely, for popular and random objects, textual data often provides sufficient statistical information, reducing the necessity for increased visual input reliance.
Table 11: More Results on POPE [29].
Dataset Setting Method LLaVA-v1.5 InstructBLIP mPLUG-Owl2
Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$
MSCOCO Random base 86.5 84.8 87.1 85.7 86.0 84.4
ours 86.8 85.1 87.9 86.8 87.9 87.1
Popular base 85.5 83.8 84.2 83.6 84.6 83.2
ours 85.6 84.0 85.0 84.3 86.4 85.7
A-OKVQA Random base 88.0 87.6 88.5 88.5 86.5 85.7
ours 88.1 87.4 88.8 88.8 88.4 88.1
Popular base 85.5 85.1 81.9 83.1 82.4 82.2
ours 85.5 85.1 82.3 83.4 85.1 85.3
Adversarial base 79.1 79.9 74.8 77.9 74.7 76.9
ours 79.5 80.1 75.3 78.2 78.2 79.9
GQA Random base 88.9 88.2 87.2 87.1 85.2 84.0
ours 88.9 88.2 87.2 87.2 86.1 85.0
Popular base 84.1 84.1 78.6 80.4 78.7 78.5
ours 84.2 84.1 78.8 80.4 81.0 80.5
Adversarial base 80.8 81.3 75.9 78.4 76.4 76.8
ours 81.1 81.6 76.1 78.5 79.2 79.1
Appendix I Different Sampling Strategies
Tab. 12 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD [26], where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
Table 12: Ablation Study on Sampling Strategies on POPE MSCOCO Adversarial Dataset.
strategy Method LLaVA-v1.5 InstructBLIP mPLUG-Owl2
Acc F1 Acc F1 Acc F1
non-greedy base $79.0_{\pm 0.51}$ $81.1_{\pm 0.53}$ $71.6_{\pm 0.49}$ $74.7_{\pm 0.46}$ $71.5_{\pm 0.30}$ $76.6_{\pm 0.28}$
ours 82.3±0.27 81.1±0.31 82.2±0.29 81.8±0.25 83.2±0.27 82.9±0.26
greedy base 79.4 81.6 79.8 81.4 72.5 77.5
ours 83.5 82.1 82.5 82.1 84.2 83.7
Appendix J Different Model Size
We evaluate our method on different model sizes, 7B and 13B, for LLaVA-v1.5 and InstructBLIP, as shown in Tab. 13. The results indicate consistent improvements across various model sizes. In each model series, the smaller model gets a larger performance boost. With our method, we can achieve high accuracy and detection rates with a smaller 7B model, outperforming a 13B model at its original performance level.
Table 13: Ablation Study on Model Size on LLaVA-QA90 Dataset.
Mothod LLaVA-v1.5 InstructBLIP
7B 13B 7B 13B
Acc Det Acc Det Acc Det Acc Det
base 3.23 3.54 4.78 4.2 3.84 4.07 5.67 4.88
ours 6.20 5.14 7.36 6.5 6.28 4.77 6.42 5.99
Appendix K Gradient Computation Details and Efficiency Ablation
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and facilitating straightforward reproduction. For comparison, we evaluate Integrated Gradients (IG) [11, 34, 21] using the SIG [11] implementation; Table 14 presents this ablation. In the table, “IG” denotes the SIG-based results, while “direct” refers to our torch.autograd.grad approach. Both methods yield comparable accuracy and F1-score, but the direct-gradient variant is substantially more efficient.
Table 14: Ablation Study on Gradient Methods on the POPE MSCOCO Adversarial Dataset
Methods Accuracy F1 Average Speed(ms)
MPLUG-Owl2 IG [11] 83.4 82.9 20335
direct 84.2 83.7 385
Appendix L Norm Selection for Token Influence
[…]
Table 15: Norm Strategies on POPE MSCOCO Adversarial Dataset.
Norm LLaVA-v1.5 InstructBLIP mPLUG-Owl2
Acc F1 Acc F1 Acc F1
L1 83.5 82.1 82.5 82.1 84.2 83.7
L2 83.2 81.9 79.5 79.6 83.2 82.9
L 83.4 82.0 82.1 81.8 80.8 80.6
Appendix M Hyper parameter Study
Maximum $\alpha_{m}$. To determine the optimal maximum amplification factor ($\alpha_{m}$) and understand its impact on the model performance, we conducted a search over values from 1 to 6 using the LLaVA-v1.5 model on the POPE dataset for discriminative tasks. For open-ended generation tasks on a subset of the MSCOCO dataset following [10], we observed garbled text when $\alpha_{m}$ was set to 5; therefore, we limited our search to values from 1 to 4. As shown in Tab. 16 and Tab. 17, discriminative task on POPE is less sensitive to the value of $\alpha_{m}$. The performance on the generative task gets improved while $\alpha_{m}$ increases but later drops. Therefore, in our experiments, the optimal maximum amplification factor ($\alpha_{m}$) is set to 5 and 3 for discriminative and generation tasks, respectively.
Table 16: $\alpha_{m}$ Study For Discriminative Task On POPE [29] in MSCOCO Adversarial Setting.
Maximum $\alpha_{m}$ 1 2 3 4 5 6
Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$
LLaVA-v1.5 83.4 82.0 83.4 82.0 83.4 82.0 83.4 82.0 83.5 82.1 83.4 82.1
Table 17: Maximum $\alpha_{m}$ Study For Generation Task on the MSCOCO Subset.
Maximum $\alpha_{m}$ 1 2 3 4
$C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ LenLen $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ LenLen $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ LenLen $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ LenLen
LLaVA-v1.5 44.0 11.8 76.2 86.1 41.4 11.1 77.4 84.8 41.0 10.9 77.3 85.0 41.4 10.9 77.3 84.9
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following [10]. Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 18. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 19, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
Table 18: Early Stopping Threshold Study on the MSCOCO Subset
LLaVA-v1.5 LLaVA-v1.6 IntructBLIP mPLUG-Owl2 InternVL2
$6\%$ $7\%$ $8\%$ $9\%$ $5\%$ $7\%$ $9\%$ $11\%$ $15\%$ $20\%$ $25\%$ $30\%$ $2\%$ $2.5\%$ $3\%$ $3.5\%$ $8\%$ $10\%$ $12\%$ $14\%$
$C_{S}$ 45.6 41.0 36.6 31.4 29.0 26.0 23.0 17.8 52.6 51.4 47.4 36.0 51.8 45.0 41.2 41.0 37.0 35.2 34.6 32.9
$C_{I}$ 11.5 10.9 10.2 10.2 8.5 8.1 7.8 7.5 15.0 14.3 13.4 11.7 13.7 12.4 11.0 11.0 8.6 8.1 8.0 7.9
$R$ 79.7 77.3 75.2 70.8 68.5 63.0 58.8 53.0 75.1 74.4 72.3 68.8 77.7 74.9 73.8 73.5 65.8 65.4 65.4 64.0
$L_{en}$ 92.0 85.0 75.4 63.9 119.1 101.8 81.1 62.7 107.9 103.4 93.9 74.3 89.1 83.5 78.9 77.8 180.4 175.5 170.6 162.2
Table 19: Activation Rate of the Early Stopping on the MSCOCO Subset
Methods Activate Percentage Average Length
LLaVA-v1.5 ($7\%$) base - 85.1
ours $8.7\%$ 84.4
Appendix N Additional Implementation and Experimental Details
Noun tokens are identified using the spaCy library via its en_core_web_sm model. For experiments on the Amber dataset [44], we adopt the original data splits and evaluation metrics. In the MSCOCO [30] subset, we follow the data partitioning and evaluation protocol of Deng et al. [10], with splits available in their official repository. For the LLaVA-QA90 [32], MME [13] and POPE [29] datasets, our setup replicates that of Leng et al. [26] and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench [33], MM-Vet [49] follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
Appendix O Question Category Results on the AMBER Dataset
We report discriminative results across different question categories in Tab. 20. Our method improves performance in nearly all categories across all MLLMs. The improvement in InternVL2’s object existence is minor, likely due to its already high visual influence ratio. For LLaVA-v1.5 and mPLUG-Owl2, which have lower original visual influence ratios, our method achieves more substantial gains in existence, attribute, and state categories.
Table 20: Results on the Question Categories of Discriminative Task on AMBER Dataset.
Category Metric InstructBLIP LLaVA-v1.5 LLaVA-v1.6 mPLUG-Owl2 Intern-VL2
base ours base ours base ours base ours base ours
Existence acc 70.0 79.8 70.8 93.2 92.9 93.0 75.2 89.9 90.6 90.6
P 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
R 70.0 79.8 70.8 93.2 92.9 93.0 75.2 89.9 90.6 90.6
F1 82.3 88.7 82.9 96.4 96.3 96.3 85.8 94.6 95.0 95.0
Attribute acc 71.9 78.3 72.3 76.1 75.2 77.1 73.9 78.2 82.3 82.6
P 76.0 81.7 87.3 74.0 74.6 76.4 86.0 76.9 80.9 80.9
R 64.3 73.0 52.2 82.7 83.0 83.9 57.1 81.8 84.7 85.2
F1 69.7 77.1 65.3 78.1 78.5 80.0 68.6 79.3 82.8 83.0
State acc 73.4 76.4 68.2 73.3 78.6 75.2 70.5 77.9 81.2 81.2
P 75.1 77.1 86.2 70.3 78.6 74.7 84.9 75.5 79.1 78.7
R 70.6 75.3 43.3 82.0 78.5 82.9 49.8 83.1 84.8 85.5
F1 72.8 76.2 57.6 75.7 78.5 78.6 62.8 79.1 81.8 82.0
Number acc 65.4 80.6 75.0 80.1 80.1 80.2 77.8 76.5 82.6 83.3
P 75.4 93.1 86.9 79.1 79.2 78.6 86.0 77.0 83.0 84.0
R 45.8 66.2 59.5 82.4 81.7 84.4 66.9 77.0 82.0 82.3
F1 57.0 77.4 70.6 80.7 80.4 81.4 75.3 77.0 82.5 83.1
Action acc 79.7 83.7 83.6 82.3 81.9 80.4 84.0 84.1 88.4 88.6
P 82.5 88.5 92.9 85.9 79.4 81.2 90.9 85.9 86.5 86.8
R 75.3 77.5 72.7 87.4 86.0 88.6 75.5 85.9 90.9 91.2
F1 78.7 82.6 81.6 86.6 82.6 84.7 82.5 85.9 88.6 88.9
Relation acc 62.7 71.9 71.8 61.5 64.5 65.7 70.5 76.9 72.1 77.0
P 56.2 64.0 65.9 51.9 54.0 56.6 61.0 67.9 60.0 65.1
R 48.6 73.4 66.3 97.7 95.1 87.2 79.5 83.9 98.3 95.6
F1 52.1 68.4 66.1 67.8 68.9 68.6 69.0 75.1 74.5 77.5
Appendix P Confidence and Visual Influence
Low confidence often signals potential failure modes in base MLLMs. Here, we demonstrate that our method not only improves accuracy but also increases model confidence. It remains effective even in low-confidence regions for three reasons: 1) We aggregate token gradients at the component level (Eq. 5) rather than using individual token gradients which yields robustness against local gradient noise. 2) We adjust influence towards visual tokens which consistently reduces the hallucination likelihood; 3) Empirically, low model confidence does not correlate with noisy gradients. In our experiments, pretrained MLLMs usually maintain meaningful gradient signals even at low confidence levels. Fig. 6 shows an example where the baseline model mPLUG-Owl2 exhibits low confidence in hallucinated predictions and near-zero confidence in the initial predictions for ‘forks’ and ‘mug’. With GACD, prediction confidence increases alongside the visual influence ratio, with the minimum confidence rising to over $30\%$.
(a) mPLUGOwl2
(b) mPLUGOwl2-GACD
Figure 6: Comparison of prediction confidence with and without GACD. (a) Without GACD, mPLUGOwl2 exhibits low confidence in hallucinated predictions and near-zero confidence in the initial predictions for ‘forks’ and ‘mug’. (b) With GACD, mPLUGOwl2’s confidence increases alongside the visual influence ratio, effectively mitigating hallucinations.
Appendix Q Qualitative Example on Occluded Images
We include a qualitative example in Fig. 7, where a sedan and a building are partially occluded by a white truck, our method prevents the baseline model from hallucinating of persons and vehicles behind the occluding object. This demonstrates our method remains effective on images consisting of occlusions.
Figure 7: Example of our method applied to an occluded image.
Image caption w/o Ours: The image features a white truck parked on a city street, with graffiti covering its side. The truck is positioned near a crosswalk, and there are several other vehicles in the scene, including a car and a bus. In addition to the vehicles, there are a few people walking around the area. Image caption w Ours: The image features a large white truck parked on a city street, with graffiti covering its side. The truck is positioned near a crosswalk, and there is another vehicle visible in the background. Additionally, there are buildings in the scene, suggesting an urban setting.
Appendix R Computational Cost
In this section, we analyze the computational cost of the proposed method. Typically, gradient computation adds an overhead equivalent to 1–2 times the cost of a forward pass. As a result, our approach increases the overall computational cost by up to five times compared to the baseline while remaining significantly lower than fine-tuning-based methods. Notably, contrastive logits generation involves only a small number of visual tokens, helping to reduce the actual runtime cost.
Appendix S Influence Ratio in Image Caption
We further visualize the influence ratio in the image captioning task. Fig. 8 shows that in the baseline LLaVA-v1.5, the influence gap between previous output tokens and visual tokens widens as more tokens are generated. However, GACD effectively narrows this gap, preventing visual information from being forgotten and thereby reducing hallucinations.
Figure 8: Comparison of influence ratios across predicted tokens with and without GACD. (Left) Without GACD, the influence gap between previous output tokens and visual tokens widens as more tokens are generated. (Right) With GACD, the gap is periodically narrowed to nearly zero, mitigating this trend and reducing hallucination.
Appendix T Influence Ratio in VQA
Fig. 9 illustrates the visual influence ratio across outputs in VQA tasks, comparing baseline predictions with those obtained after applying GACD. The results confirm that text tokens dominate influence across MLLMs, including InternVL2, which exhibits a relatively high visual influence ratio. As shown in Fig. 3 of the main paper, the overall $60\%$$100\%$ visual influence ratio across the POPE dataset suggests that visual inputs predominantly determine object existence in VQA tasks. GACD enhances visual influence, effectively balancing text-visual bias. Furthermore, the visualization on InternVL2 demonstrates that the co-occurrence hallucination ‘knife’ persists despite a high visual influence. GACD successfully eliminates this co-occurrence hallucination.
Appendix U Limitations and Future Work
Our method is limited to white-box MLLMs, as it requires access to gradients. Its effectiveness depends on the baseline MLLM’s original visual influence ratio, and the importance of visual information. As a post-processing technique, our method does not involve model training. In future work, we aim to explore how insights from GACD can guide and improve training strategies for enhanced visual perception in MLLMs.
Appendix V Broader Impacts
Our method enhances the factual reliability of multi-modal language models, not only for vision–language tasks but also for modalities such as video and audio, by mitigating hallucinations at inference time. This improvement has several positive societal implications: it can make systems for visual question answering, assistive technologies for the visually impaired, and automated image captioning more dependable, thereby increasing user trust and safety; it can power educational tools that generate accurate descriptions of complex diagrams or historical media, benefiting learners and instructors; and in critical domains such as medical imaging or remote sensing, it can reduce spurious outputs and support more robust decision-making. Conversely, if deployed within surveillance or facial-recognition systems, stronger multi-modal grounding could facilitate more intrusive inferences about individuals from visual data, exacerbating privacy risks.
(a) What are the main objects on the table in the image?
(b) LLaVAv1.5
(c) LLaVAv1.5-GACD
(d) InstructBLIP
(e) InstructBLIP-GACD
(f) InternVL2
(g) InternVL2-GACD
Figure 9: Influence Ratio across Predicted Tokens in VQA: (left) Baseline predictions; (right) Predictions with GACD. GACD effectively mitigate Text-Visual GAP, balancing text-visual bias. (f) The original InternVL2 shows a dominant visual influence ratio at the hallucinated prediction ‘knife’, indicating a co-occurrence bias that remains unaddressed even with dominant visual influence. (g) GACD successfully eliminates co-occurrence hallucinations, including ‘knife’.